If average cost=Q+5+36/Q, find out the level of output for which average cost is minimum? Show,also, the slope of average cost.
1
Expert's answer
2021-11-01T12:47:47-0400

Solution:

AC = Q + 5 + 36/Q

Derive the derivative and set it to zero to determine the output.

"\frac{\partial AC} {\partial Q} = -\frac{36} {Q^{2} } + 1"


Q = 6


The level of output for which the average cost is minimum = 6

 

The slope of the average cost curve equals the marginal cost curve at its minimum point. It is upward sloping.